# Adaptive Bounding Box Uncertainties via Two-Step Conformal Prediction

### v1

@@ 0.C.2 Prior work comparison and other results
Figure 7: Comparison of our modified one-sided approaches to conformal scoring methods proposed by [1] (see subsection 0.B.5) in terms of $MPIW$ (top) and $Stretch$ (bottom). We find that our approaches are competitive with regards to both evaluation metrics.
Figure 7 displays the results for the comparison of our CP approaches to the proposed conformal methods from Andéol et al. [1] (see subsection 0.B.5). We observe that our modified one-sided approaches provide competitive performance, as measured via efficiency metrics $MPIW$ and $Stretch$, and our best approach slightly outperforms prior work.
Table 5 displays results for the conformal box procedures from section 4 across all three datasets using the Bonferroni correction rather than $\max$-rank to account for multiple testing (see subsection 4.1). In comparison to results using $\max$-rank in Figure 3, Figure 9, and Figure 10, we observe generally inferior performance. This is expressed by a higher inefficiency measured via larger $MPIW$ and a tendency to overcover beyond the target level $(1-\alpha_{B})\approx 0.9$. The effects are particularly visible for datasets COCO and Cityscapes, due to the influence of its smaller calibration set sizes.
We further ablate results of our two-step approach using ClassThr by stratifying efficiency metrics across object (mis-)classification in Figure 8 for COCO. Results confirm expected behaviour, in that the model’s higher uncertainty for more ambiguous objects – as reflected by their misclassification – tends to align with obtained larger prediction sets.
@@
Table 5: Metrics comparison of proposed conformal methods for box coordinates (see section 4) across the three datasets using the Bonferroni correction for multiple testing. Values are means over trials and the selected set of classes.
@@ 0.C.3 Predictive performance
We validate the predictive performance – as measured via average precision (AP) metrics following the COCO detection challenge [^5] https://cocodataset.org/#detection-eval – for our primary pretrained object detection model X101-FPN across datasets in Table 6. Obtained scores are similar to results reported by detectron2666https://github.com/facebookresearch/detectron2/blob/main/MODEL_ZOO.md, confirming that the model performs adequately as a base predictor on which we run our conformal procedures. Other pretrained models employed for Box-Ens perform similarly.
@@
Table 6: Average precision (AP) scores following the COCO detection challenge metrics for our primarily employed pretrained object detection model X101-FPN (see subsection 0.B.4). Results are the mean over our selected set of classes. The primary metric AP@IoU=.50:.05:.95 averages AP scores for 10 different IoU thresholds in $[0.5,0.95]$ with step size $0.05$. AP-small, AP-med and AP-large compute scores stratified across object sizes (see our adaptivity metric in subsection 6.1).
Figure 9: Top: Empirical coverage levels marginally across all objects (All) and across objects from selected classes for the three conformal bounding box methods on the Cityscapes dataset. Bottom: Coverage levels stratified by object size (Small, Medium, Large) and $MPIW$. We also visualize target coverage (Refer to caption) and the marginal coverage distribution (Refer to caption). Displayed densities are results obtained over 1000 trials. For interpretation see subsection 6.2.
Figure 10: Top: Empirical coverage levels marginally across all objects (All) and across objects from selected classes for the three conformal bounding box methods on the Cityscapes dataset. Bottom: Coverage levels stratified by object size (Small, Medium, Large) and $MPIW$. We also visualize target coverage (Refer to caption) and the marginal coverage distribution (Refer to caption). Displayed densities are results obtained over 1000 trials. For interpretation see subsection 6.2.
@@ 0.C.4 Different target coverage levels
Figure 11: We perform ablations on different target combinations of box coverage $(1-\alpha_{B})$ and label coverage $(1-\alpha_{L})$ for a two-step conformal approach using Box-Std and ClassThr on COCO (left), Cityscapes (middle) and BDD100k (right). We denote at each intersection of dashed lines (Refer to caption) the desired target combination. We also display obtained efficiencies, as measured via $MPIW$ for conformal box intervals and mean set size for conformal label sets. Results are averaged across the selected set of classes and 100 trials.
We evaluate a series of different target combinations of box coverage $(1-\alpha_{B})$ and label coverage $(1-\alpha_{L})$ for a two-step conformal approach using Box-Std and ClassThr on all the datasets in Figure 11. Specifically, we consider all combinations of $(1-\alpha_{B})\in\{0.85,0.9,0.95\}$ and $(1-\alpha_{L})\in\{0.8,0.9,0.99,1.0\}$. Note that the combination $(1-\alpha_{B})=0.9,(1-\alpha_{L})=0.99$ is the primary coverage goal throughout this work (see also subsection 5.1).
We compare the obtained empirical coverage levels against desired target levels, denoted by the intersections of dashed lines in the top figure row. Target coverage levels are mostly satisfied with regards to box coverage, albeit the target is reached more clearly as calibration set sizes increase, as we move from COCO (left) to Cityscapes (middle) and BDD100k (right). In contrast, target levels for label coverage are fully satisfied, and in fact tend to overcover. As calibration set sizes increase, this overcoverage tendency interestingly decreases. See also subsection 0.A.6 for more details on calibration set size.
We further display obtained efficienciy, as measured via $MPIW$ for conformal box intervals and mean set size for conformal label sets, in the bottom figure row. Two trends are immediately apparent: as we increase the requirement on box coverage, $MPIW$ increases; similarly, as we increase the requirement on label coverage, mean set size increases. Interestingly, the gap between permitting minimal label miscoverage and no miscoverage at all, i.e., $\alpha_{L}=0.01$ and $\alpha_{L}=0$, tends to increase as the number of calibration samples rises. This observation motivates our choice to select $\alpha_{L}=0.01$ as a practical, approximate way of ensuring total label coverage whilst minimally affecting the downstream box guarantees.
@@ 0.C.5 Additional visualisations
We display additional examples of conformal bounding box intervals produced using our two-step approach (see Figure 2) for COCO in Figure 12, for Cityscapes in Figure 13 and for BDD100k in Figure 14. We use ClassThr to produce class label prediction sets with 99% guaranteed coverage and either Box-Std, Box-Ens or Box-CQR (from left to right by image column) to produce conformal bounding box intervals with approx. 90% guaranteed coverage. We depict a range of different classes which align with our set of selected classes of particular interest relevant for urban driving scenes: person, bicycle, motorcycle, car, bus and truck. Note that for some images we only draw intervals for a filtered amount of classes for clearer visualisations. We observe that despite its ability to ensure probabilistic guarantees even under object misclassification, the obtained bounding box uncertainty intervals are reasonably tight, and could effectively be used for improved decision-making in many instances.

### v2

@@ 0.C.2 Prior work comparison and other results
Figure 7: Comparison of our modified one-sided approaches to conformal scoring functions proposed by [1] (see § 0.B.5) in terms of $MPIW$ (top) and $Stretch$ (bottom). We find that our approaches are competitive with regards to both evaluation metrics.
Fig. 7 displays results for the comparison of our bounding box methods to the proposed scoring functions from Andéol et al. [1] (see § 0.B.5). We observe that our modified one-sided approaches are competitive, as measured via efficiency metrics $MPIW$ and $Stretch$, and our best approach even slightly outperforms.
Tab. 5 displays results for the bounding box methods from § 4 across all three datasets using the Bonferroni correction rather than $\max$-rank to account for multiple testing (see § 4.1). In comparison to results using $\max$-rank (Fig. 3, Fig. 10, and Fig. 11), performance is generally inferior. This is captured by larger $MPIW$ and a tendency to overcover beyond the target level of $(1-\alpha_{B})\approx 0.9$. The effects are particularly visible for datasets COCO and Cityscapes.
We further ablate results of our two-step approach using ClassThr by stratifying efficiency metrics across object (mis-)classification in Fig. 8 for COCO. Results confirm that the model’s higher uncertainty for more ambiguous objects – as reflected by their misclassification – tends to result in larger prediction sets.
@@
Table 5: Metrics comparison of proposed conformal methods for box coordinates (see § 4) across the three datasets using the Bonferroni correction for multiple testing. Values are means over trials and the selected set of classes.
@@ 0.C.3 Predictive performance
We validate the predictive performance – as measured via average precision (AP) metrics following the COCO detection challenge  [^7] See https://cocodataset.org/#detection-eval. – for our primary pretrained object detection model X101-FPN across datasets in Tab. 6. Obtained scores are similar to results reported by detectron2  [^8] https://github.com/facebookresearch/detectron2/blob/main/MODEL_ZOO.md, confirming that the model performs adequately as a base predictor on which we run our conformal procedures. Other pretrained models employed for Box-Ens perform similarly.
@@
Table 6: Average precision (AP) scores following the COCO detection challenge metrics for our primarily employed pretrained object detection model X101-FPN (see § 0.B.4). Results are the mean over our selected set of classes. The primary metric AP@IoU=.50:.05:.95 averages AP scores for 10 different IoU thresholds in $[0.5,0.95]$ with step size $0.05$. AP-small, AP-med and AP-large compute scores stratified across object sizes (see our adaptivity metric in § 6.1).
@@ 0.C.4 Different target coverage levels
Figure 9: We ablate different combinations of box coverage $(1-\alpha_{B})$ and label coverage $(1-\alpha_{L})$ for a two-step conformal approach using Box-Std and ClassThr on COCO (left), Cityscapes (middle) and BDD100k (right). We denote at each intersection of dashed lines (Refer to caption) the desired target combination. We also display obtained efficiencies, as measured via $MPIW$ for conformal box intervals and mean set size for conformal label sets. Results are averaged across the selected set of classes and 100 trials.
We evaluate a series of different target combinations of box coverage $(1-\alpha_{B})$ and label coverage $(1-\alpha_{L})$ for our two-step conformal approach using Box-Std and ClassThr on all the datasets in Fig. 9. Specifically, we consider all combinations of $(1-\alpha_{B})\in\{0.85,0.9,0.95\}$ and $(1-\alpha_{L})\in\{0.8,0.9,0.99,1.0\}$. Note that the combination $(1-\alpha_{B})=0.9,(1-\alpha_{L})=0.99$ is the primary coverage goal throughout this work (see also § 5.1). We compare the obtained empirical coverage levels against desired target levels, denoted by the intersections of dashed lines in the top figure row. Target coverage levels are mostly satisfied with regards to box coverage, albeit the target is reached more clearly as calibration set sizes increase, as we move from COCO (left) to Cityscapes (middle) and BDD100k (right). In contrast, target levels for label coverage are fully satisfied, and in fact tend to overcover. As calibration set sizes increase, this overcoverage tendency decreases. See also § 0.A.6 for more details on calibration set size.
We further display obtained efficiency, as measured via $MPIW$ for box intervals and mean set size for label sets (bottom row). Two trends are immediately apparent: as we increase the requirement on box coverage, $MPIW$ increases; similarly, as we increase the requirement on label coverage, mean set size increases. Interestingly, the gap between permitting minimal label miscoverage and no miscoverage at all, i.e., $\alpha_{L}=0.01$ and $\alpha_{L}=0$, tends to increase as the number of calibration samples rises. This observation motivates our choice to select $\alpha_{L}=0.01$ as a practical approximation ensuring total label coverage whilst minimally affecting the downstream box guarantees.
Figure 10: Top: Empirical coverage levels marginally across all objects (All) and across objects from selected classes for the three conformal bounding box methods on the Cityscapes dataset. Bottom: Coverage levels stratified by object size (Small, Medium, Large) and $MPIW$. We also visualize target coverage (Refer to caption) and the marginal coverage distribution (Refer to caption). Displayed densities are results obtained over 1000 trials. For interpretation see § 6.2.
Figure 11: Top: Empirical coverage levels marginally across all objects (All) and across objects from selected classes for the three conformal bounding box methods on the Cityscapes dataset. Bottom: Coverage levels stratified by object size (Small, Medium, Large) and $MPIW$. We also visualize target coverage (Refer to caption) and the marginal coverage distribution (Refer to caption). Displayed densities are results obtained over 1000 trials. For interpretation see § 6.2.
@@ 0.C.5 Additional visualisations
We display additional examples of conformal bounding box intervals produced using our two-step approach for COCO in Fig. 12, for Cityscapes in Fig. 13 and for BDD100k in Fig. 14. We use ClassThr to produce class label prediction sets with 99% guaranteed coverage and either Box-Std, Box-Ens or Box-CQR (from left to right by image column) to produce box intervals with $\sim$90guaranteed coverage. We depict a range of different classes which align with our particular interest in urban driving scenes: person, bicycle, motorcycle, car, bus and truck. Note that for some images we only visualize a filtered amount of classes for clarity. We observe that despite its ability to ensure probabilistic guarantees even under object misclassification, the obtained bounding box intervals are reasonably tight, and could effectively be used for decision-making.